# Minimal Subsampled Rank-1 Lattices for Multivariate Approximation with Optimal Convergence Rate

### v2

@@ 5 Application to weighted Korobov spaces
6 Implementation and computational cost
@@
$\frac{\|\bm{x}^{(i)}-\bm{\tilde{x}}\|_{2}}{\sqrt{\kappa_{2}(\bm{A})}\|\bm{\tilde{x}}\|_{2}}\leq\sqrt{\frac{(\bm{x}^{(i)}-\bm{\tilde{x}})^{\ast}\bm{A}\,(\bm{x}^{(i)}-\bm{\tilde{x}})}{(\bm{\tilde{x}})^{\ast}\bm{A}\,\bm{\tilde{x}}}}\leq 2\Big{(}\frac{\sqrt{\kappa_{2}(\bm{A})}-1}{\sqrt{\kappa_{2}(\bm{A})}+1}\Big{)}^{i}\leq 2\exp\Big{(}-\frac{2i}{\sqrt{\kappa_{2}(\bm{A})}}\Big{)}\,.$
@@
The LSQR method is analytically equivalent to applying CG to the normal equation but in a numerically more stable manner, cf. [PS82]. Thus the iteration bound in Theorem 6.1 holds for LSQR with the minimal and maximal eigenvalue replaced by the squared singular numbers.
6.1 Fast least squares approximation with subsampled lattices
@@
be defined in (4.1). The approximation is computed once the Fourier coefficients $\bm{\widehat{g}}_{\mathcal{B}}=(\widehat{g}_{\bm{h}})_{\bm{h}\in B}$ representing the approximation are obtained. Essential is the set of frequencies $\mathcal{B}$. Because of the downward-closed structure, this can be set up incrementally in dimension. For the step from dimension $j$ to $j+1$, we have frequencies supported in the first $j$ dimensions and for every one of them we seek all $h_{j+1}$ such that $r_{d,\alpha,\bm{\gamma}}([h_{1},\dots,h_{j},h_{j+1},$ $0,\dots,0])\leq M$. This has a computational cost of $\mathcal{O}(d\,|\mathcal{B}|)$. We then compute the coefficients by solving the system of equations in Lemma 4.1 via LSQR.
By (4.6) in Lemma 4.4, we have, with $|J|$ sufficiently large and high probability, that $\kappa_{2}(\bm{L}_{J,\mathcal{B}}^{\ast}\,\bm{L}_{J,\mathcal{B}})=\lambda_{\max}(\bm{L}_{J,\mathcal{B}}^{\ast}\,\bm{L}_{J,\mathcal{B}})/\lambda_{\min}(\bm{L}_{J,\mathcal{B}}^{\ast}\,\bm{L}_{J,\mathcal{B}})\leq 3$. Fixing the desired accuracy to machine precision $\texttt{eps}=10^{-16}$, we obtain by Theorem 6.1 a fixed number of iterations for LSQR. Each iteration uses one matrix-vector product with $\bm{L}_{J,\mathcal{B}}$ and one with $\bm{L}_{J,\mathcal{B}}^{\ast}$. For the full lattice $J=\{1,\dots,n\}$ this matrix-vector product simplifies to a one-dimensional FFT with a computational cost of
@@
$\bm{L}_{J,\mathcal{B}}=\bm{P}\bm{L}_{\mathcal{B}}\quad\text{with}\quad\bm{P}=\left(\begin{matrix}\\ \\ \\ \\ \end{matrix}\right.\begin{matrix}\\ 1&0&0&\dots&0&0&0\\ 0&0&1&\dots&0&0&0\\ &\vdots&&&&\vdots&\\ 0&0&0&\dots&0&1&0\\ {\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}j_{1}}&&{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}j_{2}}&&&{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}j_{|J|}}\end{matrix}\left.\begin{matrix}\\ \\ \\ \\ \end{matrix}\right){\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\begin{matrix}1\\ 2\\ \vdots\\ |J|\end{matrix}}\in\{0,1\}^{|J|\times n}\,,$ (6.1)
@@
With $|J|\sim\sqrt{n}\log n$ and $|\mathcal{B}|\sim\sqrt{n}$ as in Theorem 5.3 we obtain the computational cost in the upper two rows of Table 1 when using the LFFT compared to the naive matrix-vector product. The iteration count does not appear here as it is bounded by a general constant independent of the number of points.
@@
Table 1: Computational cost for the least squares approximation $S_{\mathcal{B}}^{\bm{X}_{J}}$ and the kernel method $A_{\text{ker}}^{\bm{X}_{J}}$ with a subsampled lattice of size $|J|=\lceil\sqrt{n}\,\log n\rceil$ utilizing the FFT compared to the naive matrix-vector product. Here $r$ is the number of iterations depending on the condition number of the kernel matrix.
Note, when using the full lattice an iterative solver is not needed, since we have by the reconstructing property $(\bm{L}_{\mathcal{B}}^{\ast}\,\bm{L}_{\mathcal{B}}^{\ast})^{-1}=(1/n)\,\bm{I}$. This can be also viewed as LSQR converging with one iteration because $\kappa_{2}(\bm{L}_{\mathcal{B}}^{\ast}\,\bm{L}_{\mathcal{B}})=1$ in Theorem 6.1. Thus, using the full lattice has computational cost $\mathcal{O}(n\log n)$ as well.
6.2 Fast kernel method with subsampled lattices
We further comment on the kernel method, which we introduced in Section 2 for the full lattice. Since $H_{d,\alpha,\bm{\gamma}}$ is a reproducing kernel Hilbert space, it has an associated reproducing kernel $K(\cdot,\cdot)=K_{d,\alpha,\bm{\gamma}}(\cdot,\cdot)$. For integer smoothness $\alpha\in\mathds{N}$, a simple closed form is known:
$K(\bm{x},\bm{y})=\!\!\sum_{\mathfrak{u}\subseteq\{1,\dots,d\}}\!\!\gamma_{\mathfrak{u}}\prod_{j\in\mathfrak{u}}\eta_{\alpha}(x_{j},x_{j}^{\prime})\quad\text{with}\quad\eta_{\alpha}(x,x^{\prime})=\frac{(2\pi)^{2\alpha}B_{2\alpha}((x-x^{\prime})\bmod{}1)}{(-1)^{\alpha+1}(2\alpha)!}\,,$
@@ 6.2 Fast kernel method with subsampled lattices
$A_{\text{ker}}^{\bm{X}_{J}}f\coloneqq\sum_{k\in J}a_{k}\,K(\cdot,\tfrac{k}{n}\bm{z}\bmod{}1)\quad\text{with}\quad\bm{a}=(a_{k})_{k\in J}=\bm{K}_{J}^{-1}\bm{f}_{J}$ (6.2)
@@
As discussed in Section 2, the kernel method is optimal in the worst-case setting for any given set of points. In particular, the bound from Theorem 5.3 applies as well, since $e^{\mathrm{wor-app}}(A_{\text{ker}}^{\bm{X}_{J}})\leq e^{\mathrm{wor-app}}(S_{\mathcal{B}}^{\bm{X}_{J}})$. Similar to $\bm{L}_{J,\mathcal{B}}$, a fast multiplication with the kernel matrix for the full lattice is known, since $\bm{K}_{\{0,\dots,n-1\}}$ is a circulant matrix, cf. [KKKNS21, Table 1], with a computational cost of
@@
In contrast to the least squares matrix $\bm{L}_{J,\mathcal{B}}$, the condition number of the kernel matrix $\bm{K}_{J}$ grows in the number of points, resulting in the need for many iterations or the necessity of preconditioning, cf. [W04, Section 12.2]. Thus, the number of necessary CG iterations cannot be bounded by a general constant according to Theorem 6.1.
With $|J|\sim\sqrt{n}\log n$ as in Theorem 5.3, we obtain the computational cost in the lower two rows of Table 1 when using the FFT compared to the naive matrix-vector product when using $r$ iterations.
6.3 Comparison to full-lattice methods
From Theorem 5.3 we know that subsampling allows for a better sampling complexity. However, picking the best method in regards to computational cost also depends on whether functions evaluations are expensive. To discuss this balancing, we split the computations into five steps and address their individual computational cost. This gives the following for the least squares approximation and the kernel method with the full and subsampled lattice, which is summarized in Table 2.
@@ 6.3 Comparison to full-lattice methods
    Point construction. All methods use the CBC construction from [GS25] with different cost depending on the type of weights as shown in the first row of Table 2. The additional cost for creating the subsampled indices is negligible.
@@
    The sampling of the function to approximate. When $f_{\text{cost}}$ models the cost of a single function evaluation, this has to be multiplied by the number of samples used by the approximation. For the full lattice these are $n$ points and for the subsampled lattice we have $|J|\sim\sqrt{n}\log n$ points, cf. Theorem 5.3.
@@
    For the least squares approximation this involves the construction of the frequency index set $\mathcal{B}$ and the computation of the inner products with the generating vector $\bm{z}$. As discussed in Section 6.1 this can be done incrementally in dimension with a cost of $d\,|\mathcal{B}|$, where $|\mathcal{B}|\sim\sqrt{n}$ in our case.
    For the kernel method the entries of the kernel matrix $\bm{K}_{J}$ have to be computed, where the cost of one single kernel evaluation depends on the weight parameters, cf. [KKKNS21, Table 1]. Because we have $K(\frac{k}{n}\bm{z}\bmod{}1,\frac{k^{\prime}}{n}\bm{z}\bmod{}1)=K(\frac{k-k^{\prime}}{n}\bm{z}\bmod{}1,\bm{0})$, the number of different values equals |$|\{(k-k^{\prime})\bmod n:k,k^{\prime}\in J\}|$. For the full lattice this evaluates to $n$. For the randomly subsampled lattice with $|J|\sim\sqrt{n}\log n$ we might have |$|\{(k-k^{\prime})\bmod n:k,k^{\prime}\in J\}|=n$ as well.
@@
    The computation of the coefficients. At this step the approximation is obtained from the sampled function values using the function-independent setup and encoded in a respective representation. The least squares approximation is described by its Fourier coefficients $[\widehat{g}_{\bm{h}}]_{\bm{h}\in\mathcal{B}}$ and the kernel method by the coefficients $[a_{k}]_{k\in J}$ (where $J=\{0,\dots,n-1\}$ for the full lattice). These coefficients are defined by the solution of the systems of equations in Lemma 4.1 and (2.7), respectively. For the full lattice the involved inverse matrices are known in closed form as discussed in Section 6.1 and Section 2, which results in a computational cost of an FFT for the computation of the coefficients.
    For the subsampled methods we use the iterative schemes described in Section 6.1 and Section 6.2. In case of the least squares method, the number of iterations is bounded and by a general constant. For kernel method more iterations are needed when the number of points increases, as the kernel matrix is worse conditioned for an increasing number of points. In Table 2 the number of iterations for the kernel method is denoted by $r$.
@@
    The evaluation of the approximation. This describes the cost of evaluating the approximation in a single unseen data point $\bm{x}\in\mathds{T}^{d}$. For the least squares approximation this involves evaluating $(S_{\mathcal{B}}^{\bm{X}_{J}}f)(\bm{x})=\sum_{\bm{h}\in\mathcal{B}}\widehat{g}_{\bm{h}}\exp(2\pi\mathrm{i}\langle\bm{h},\bm{x}\rangle)$ with a cost of $d\,|\mathcal{B}|$, with $|\mathcal{B}|\sim\sqrt{n}$ in our case.
@@
Note that for a given $n$ all methods have the $L_{2}$ worst-case error proportional to $4]{S_{n}(\bm{z})}$, cf. Theorems 2.1 and 5.3.
@@
Table 2: Computational cost for the least squares approximation $S_{\mathcal{B}}^{\bm{X}_{J}}$ and the kernel method $A_{\text{ker}}^{\bm{X}_{J}}$ with a full and subsampled lattice of size $|J|=\lceil\sqrt{n}\,\log n\rceil$. Here $r$ is the number of iterations depending on the condition number of the kernel matrix and $f_{\text{cost}}$ is the cost of a single function evaluation. We have $(i,j)=(1,0)$ for product weights, $(i,j)=(2,0)$ for POD weights, and $(i,j)=(2,2)$ for SPOD weights.

### v3

@@ 5. Application to weighted Korobov spaces
The following Corollary shows that the asymptotic rates in Theorem˜5.3 are optimal.
Corollary 5.4.
Under the setting of Theorem˜5.3 with prime $n$, we have the asymptotic rates
(5.7) $e^{\mathrm{wor-app}}(S_{\mathcal{B}}^{\bm{X}_{J}})\lesssim|J|^{-\alpha+\varepsilon}\quad\text{with}\quad|J|^{2\sqrt{1-\varepsilon/\alpha}}\lesssim n\lesssim|J|^{2/\sqrt{1-\varepsilon/\alpha}}$
for all $0<\varepsilon<\alpha$. This is optimal in that any method utilizing $|J|$ information cannot have a better rate than $|J|^{-\alpha}$, and the rate given implies an initial lattice size of $n\gtrsim|J|^{2-2\varepsilon/\alpha}$.
Proof.
Rall that (5.4) and (5.6) in Theorem˜5.3 hold for all parameter $\lambda$ satisfying $1/(2\alpha)<\lambda<1$. We reparametrize in terms of arbitrary $0<\varepsilon<\alpha$ by choosing $\lambda$ such that
$\alpha-\frac{1}{4\alpha\lambda^{2}}<\varepsilon,$
which is possible since $\lambda>1/(2\alpha)$ implies that the left-hand side can be made arbitrarily small. Thus from (5.4) we have
$e^{\mathrm{wor-app}}(S_{\mathcal{B}}^{\bm{X}_{J}})\lesssim\Big(\frac{\log|J|}{|J|}\Big)^{\frac{1}{4\alpha\lambda^{2}}}=\Big(\frac{\log|J|}{|J|}\Big)^{\alpha-(\alpha-\frac{1}{4\alpha\lambda^{2}})}\lesssim|J|^{-\alpha+\varepsilon}\,.$
Moreover, from (5.6) we have
$n\gtrsim\Big(\frac{|J|}{\log|J|}\Big)^{\frac{1}{\alpha\lambda}}=\Big(\frac{|J|}{\log|J|}\Big)^{2\sqrt{1-(\alpha-\frac{1}{4\alpha\lambda^{2}})/\alpha}}\gtrsim|J|^{2\sqrt{1-\varepsilon/\alpha}}$
and, using $\varphi(n)=n-1$ for $n$ prime,
$n\lesssim\varphi(n)\lesssim\Big(\frac{|J|}{\log|J|}\Big)^{4\alpha\lambda}=\Big(\frac{|J|}{\log|J|}\Big)^{\frac{2}{\sqrt{1-(\alpha-\frac{1}{4\alpha\lambda^{2}})/\alpha}}}\lesssim|J|^{\frac{2}{\sqrt{1-\varepsilon/\alpha}}}.$
This proves the assertion (5.7).
The optimality of the worst-case error follows from bounds on the linear width in [DTU18, Theorem 4.45]. The optimality for the initial lattice size follows from [BKUV17], since an algorithm using a subsampled lattice is inherently still using the lattice of size $n$, and achieving the rate $|J|^{-\alpha+\varepsilon}$ implies $n^{-\alpha/2}\lesssim|J|^{-\alpha+\varepsilon}$, yielding the lower bound. ∎
We remark that the $\varepsilon$ in Corollary˜5.4 incorporates both the gap that naturally arises when approximating in Korobov spaces and the additional logarithmic gap stemming from the random nature of our method. In contrast, Theorem˜5.3 clearly distinguishes between these two effects.
6. Implementation and computational cost
@@
$\frac{\|\bm{x}^{(i)}-\bm{\tilde{x}}\|_{2}}{\sqrt{\kappa_{2}(\bm{A})}\|\bm{\tilde{x}}\|_{2}}\leq\!\sqrt{\frac{(\bm{x}^{(i)}-\bm{\tilde{x}})^{\ast}\bm{A}\,(\bm{x}^{(i)}-\bm{\tilde{x}})}{(\bm{\tilde{x}})^{\ast}\bm{A}\,\bm{\tilde{x}}}}\!\leq 2\Big(\frac{\sqrt{\kappa_{2}(\bm{A})}-1}{\sqrt{\kappa_{2}(\bm{A})}+1}\Big)^{i}\!\leq 2\exp\Big(\!\frac{-2i}{\sqrt{\kappa_{2}(\bm{A})}}\Big).$
@@
The LSQR method is analytically equivalent to applying CG to the normal equation but in a numerically more stable manner, cf. [PS82]. Thus the iteration bound in Theorem˜6.1 holds for LSQR with the minimal and maximal eigenvalue replaced by the squared singular numbers.
6.1. Fast least squares approximation with subsampled lattices
@@
be defined in (4.1). The approximation is computed once the Fourier coefficients $\bm{\widehat{g}}_{\mathcal{B}}=(\widehat{g}_{\bm{h}})_{\bm{h}\in B}$ representing the approximation are obtained. Essential is the set of frequencies $\mathcal{B}$. Because of the downward-closed structure, this can be set up incrementally in dimension. For the step from dimension $j$ to $j+1$, we have frequencies supported in the first $j$ dimensions and for every one of them we seek all $h_{j+1}$ such that $r_{d,\alpha,\bm{\gamma}}([h_{1},\dots,h_{j},h_{j+1},$ $0,\dots,0])\leq M$. This has a computational cost of $\mathcal{O}(d\,|\mathcal{B}|)$. We then compute the coefficients by solving the system of equations in Lemma˜4.1 via LSQR.
By (4.6) in Lemma˜4.4, we have, with $|J|$ sufficiently large and high probability, that $\kappa_{2}(\bm{L}_{J,\mathcal{B}}^{\ast}\,\bm{L}_{J,\mathcal{B}})=\lambda_{\max}(\bm{L}_{J,\mathcal{B}}^{\ast}\,\bm{L}_{J,\mathcal{B}})/\lambda_{\min}(\bm{L}_{J,\mathcal{B}}^{\ast}\,\bm{L}_{J,\mathcal{B}})\leq 3$. Fixing the desired accuracy to machine precision $\texttt{eps}=10^{-16}$, we obtain by Theorem˜6.1 a fixed number of iterations for LSQR. Each iteration uses one matrix-vector product with $\bm{L}_{J,\mathcal{B}}$ and one with $\bm{L}_{J,\mathcal{B}}^{\ast}$. For the full lattice $J=\{1,\dots,n\}$ this matrix-vector product simplifies to a one-dimensional FFT with a computational cost of
@@
(6.1) $\bm{L}_{J,\mathcal{B}}=\bm{P}\bm{L}_{\mathcal{B}}\quad\text{with}\quad\bm{P}=\left(\begin{matrix}\\ \\ \\ \\ \end{matrix}\right.\begin{matrix}\\ 1&0&0&\dots&0&0&0\\ 0&0&1&\dots&0&0&0\\ &\vdots&&&&\vdots&\\ 0&0&0&\dots&0&1&0\\ {\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}j_{1}}&&{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}j_{2}}&&&{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}j_{|J|}}\end{matrix}\left.\begin{matrix}\\ \\ \\ \\ \end{matrix}\right){\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}\begin{matrix}1\\ 2\\ \vdots\\ |J|\end{matrix}}\in\{0,1\}^{|J|\times n}\,$,
@@
With $|J|\sim\sqrt{n}\log n$ and $|\mathcal{B}|\sim\sqrt{n}$ as in Theorem˜5.3 we obtain the computational cost in the upper two rows of Table˜1 when using the LFFT compared to the naive matrix-vector product. The iteration count does not appear here as it is bounded by a general constant independent of the number of points.
@@
Table 1. Computational cost for the least squares approximation $S_{\mathcal{B}}^{\bm{X}_{J}}$ and the kernel method $A_{\text{ker}}^{\bm{X}_{J}}$ with a subsampled lattice of size $|J|=\lceil\sqrt{n}\,\log n\rceil$ utilizing the FFT compared to the naive matrix-vector product. Here $r$ is the number of iterations depending on the condition number of the kernel matrix.
Note, when using the full lattice an iterative solver is not needed, since we have by the reconstructing property $(\bm{L}_{\mathcal{B}}^{\ast}\,\bm{L}_{\mathcal{B}}^{\ast})^{-1}=(1/n)\,\bm{I}$. This can be also viewed as LSQR converging with one iteration because $\kappa_{2}(\bm{L}_{\mathcal{B}}^{\ast}\,\bm{L}_{\mathcal{B}})=1$ in Theorem˜6.1. Thus, using the full lattice has computational cost $\mathcal{O}(n\log n)$ as well.
6.2. Fast kernel method with subsampled lattices
We further comment on the kernel method, which we introduced in Section˜2 for the full lattice. Since $H_{d,\alpha,\bm{\gamma}}$ is a reproducing kernel Hilbert space, it has an associated reproducing kernel $K(\cdot,\cdot)=K_{d,\alpha,\bm{\gamma}}(\cdot,\cdot)$. For integer smoothness $\alpha\in\mathds{N}$, a simple closed form is known:
$K(\bm{x},\bm{y})=\!\!\sum_{\mathfrak{u}\subseteq\{1,\dots,d\}}\!\!\gamma_{\mathfrak{u}}\prod_{j\in\mathfrak{u}}\eta_{\alpha}(x_{j},x_{j}^{\prime})\;\;\text{with}\;\;\eta_{\alpha}(x,x^{\prime})=\frac{(2\pi)^{2\alpha}B_{2\alpha}((x-x^{\prime})\bmod{}1)}{(-1)^{\alpha+1}(2\alpha)!}\,$,
@@ 6.2. Fast kernel method with subsampled lattices
(6.2) $A_{\text{ker}}^{\bm{X}_{J}}f\coloneqq\sum_{k\in J}a_{k}\,K(\cdot,\tfrac{k}{n}\bm{z}\bmod{}1)\quad\text{with}\quad\bm{a}=(a_{k})_{k\in J}=\bm{K}_{J}^{-1}\bm{f}_{J}\,$,
@@
As discussed in Section˜2, the kernel method is optimal in the worst-case setting for any given set of points. In particular, the bound from Theorem˜5.3 applies as well, since $e^{\mathrm{wor-app}}(A_{\text{ker}}^{\bm{X}_{J}})\leq e^{\mathrm{wor-app}}(S_{\mathcal{B}}^{\bm{X}_{J}})$. Similar to $\bm{L}_{J,\mathcal{B}}$, a fast multiplication with the kernel matrix for the full lattice is known, since $\bm{K}_{\{0,\dots,n-1\}}$ is a circulant matrix, cf. [KKKNS21, Table 1], with a computational cost of
@@
In contrast to the least squares matrix $\bm{L}_{J,\mathcal{B}}$, the condition number of the kernel matrix $\bm{K}_{J}$ grows in the number of points, resulting in the need for many iterations or the necessity of preconditioning, cf. [W04, Section 12.2]. Thus, the number of necessary CG iterations cannot be bounded by a general constant according to Theorem˜6.1.
With $|J|\sim\sqrt{n}\log n$ as in Theorem˜5.3, we obtain the computational cost in the lower two rows of Table˜1 when using the FFT compared to the naive matrix-vector product when using $r$ iterations.
6.3. Comparison to full-lattice methods
From Theorem˜5.3 we know that subsampling allows for a better sampling complexity. However, picking the best method in regards to computational cost also depends on whether functions evaluations are expensive. To discuss this balancing, we split the computations into five steps and address their individual computational cost. This gives the following for the least squares approximation and the kernel method with the full and subsampled lattice, which is summarized in Table˜2.
@@ 6.3. Comparison to full-lattice methods
    Point construction. All methods use the CBC construction from [GS25] with different cost depending on the type of weights as shown in the first row of Table˜2. The additional cost for creating the subsampled indices is negligible.
@@
    The sampling of the function to approximate. When $f_{\text{cost}}$ models the cost of a single function evaluation, this has to be multiplied by the number of samples used by the approximation. For the full lattice these are $n$ points and for the subsampled lattice we have $|J|\sim\sqrt{n}\log n$ points, cf. Theorem˜5.3.
@@
    For the least squares approximation this involves the construction of the frequency index set $\mathcal{B}$ and the computation of the inner products with the generating vector $\bm{z}$. As discussed in Section˜6.1 this can be done incrementally in dimension with a cost of $d\,|\mathcal{B}|$, where $|\mathcal{B}|\sim\sqrt{n}$ in our case.
    For the kernel method the entries of the kernel matrix $\bm{K}_{J}$ have to be computed, where the cost of one single kernel evaluation depends on the weight parameters, cf. [KKKNS21, Table 1]. Because we have $K(\frac{k}{n}\bm{z}\bmod{}1,\frac{k^{\prime}}{n}\bm{z}\bmod{}1)=K(\frac{k-k^{\prime}}{n}\bm{z}\bmod{}1,\bm{0})$, the number of different values equals $|\{(k-k^{\prime})\bmod n:k,k^{\prime}\in J\}|$. For the full lattice this evaluates to $n$. For the randomly subsampled lattice with $|J|\sim\sqrt{n}\log n$ we might have $|\{(k-k^{\prime})\bmod n:k,k^{\prime}\in J\}|=n$ as well.
@@
    The computation of the coefficients. At this step the approximation is obtained from the sampled function values using the function-independent setup and encoded in a respective representation. The least squares approximation is described by its Fourier coefficients $\widehat{g}_{\bm{h}}]_{\bm{h}\in\mathcal{B}}$ and the kernel method by the coefficients $a_{k}]_{k\in J}$ (where $J=\{0,\dots,n-1\}$ for the full lattice). These coefficients are defined by the solution of the systems of equations in Lemma˜4.1 and (2.7), respectively. For the full lattice the involved inverse matrices are known in closed form as discussed in Section˜6.1 and Section˜2, which results in a computational cost of an FFT for the computation of the coefficients. T least squares matrix $\bm{L}_{\mathcal{B}}$ for a full lattice with reconstructing property is orthogonal and therefore has condition number $\kappa_{2}(\bm{L}_{\mathcal{B}})=1$. In contrast to that, the kernel matrix $\bm{K}$ has a bigger condition number, which grows with the smoothness $\alpha$ and number of points $n$, cf. [abithesis, Chapter 6]. This might require special attention when computing the kernel approximation.
    For the subsampled methods we use the iterative schemes described in Section˜6.1 and Section˜6.2. In case of the least squares method, the number of iterations is bounded and by a general constant. For kernel method more iterations are needed when the number of points increases, as the kernel matrix is worse conditioned for an increasing number of points. In Table˜2 the number of iterations for the kernel method is denoted by $r$.
@@
    The evaluation of the approximation. This describes the cost of evaluating the approximation in a single unseen data point $\bm{x}\in\mathds{T}^{d}$. For the least squares approximation this involves evaluating the sum $(S_{\mathcal{B}}^{\bm{X}_{J}}f)(\bm{x})=\sum_{\bm{h}\in\mathcal{B}}\widehat{g}_{\bm{h}}\exp(2\pi\mathrm{i}\langle\bm{h},\bm{x}\rangle)$ with a cost of $d\,|\mathcal{B}|$, with $|\mathcal{B}|\sim\sqrt{n}$ in our case.
@@
Note that for a given $n$ all methods have the $L_{2}$ worst-case error proportional to $\sqrt[4]{S_{n}(\bm{z})}$, cf. Theorems˜2.1 and 5.3.
@@
Table 2. Computational cost for the least squares approximation $S_{\mathcal{B}}^{\bm{X}_{J}}$ and the kernel method $A_{\text{ker}}^{\bm{X}_{J}}$ with a full and subsampled lattice of size $|J|=\lceil\sqrt{n}\,\log n\rceil$. Here $r$ is the number of iterations depending on the condition number of the kernel matrix and $f_{\text{cost}}$ is the cost of a single function evaluation. We have $(i,j)=(1,0)$ for product weights, $(i,j)=(2,0)$ for POD weights, and $(i,j)=(2,2)$ for SPOD weights.